OpenStudy (anonymous):

what is the value of d? 2d – 4 = 4

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

d=4

OpenStudy (anonymous):

might be 4-0=4

OpenStudy (anonymous):

d = 4. 2 x 4 - 4 = 4

OpenStudy (anonymous):

which repesent 2d

OpenStudy (anonymous):

d=4 because 2x4=8-4=4 so, d=4 is correct

OpenStudy (anonymous):

thanks